OpenStudy (anonymous):

Select the factors of x2 - 16x + 64

OpenStudy (anonymous):

a = 1 b = -16 c = 64

OpenStudy (anonymous):

Ok thankz :)

OpenStudy (anonymous):

That wasn't answer

OpenStudy (anonymous):

factor is the number next to x^2 ,x and the absolute term

OpenStudy (anonymous):

(4x - 3)(x - 3) (4x + 3)(x + 3) (2x - 3)(2x - 3) (2x + 3)(2x + 3)

OpenStudy (anonymous):

Those are my choices

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you have to apply the quadratic formula...

OpenStudy (tkhunny):

Do not apply the Quadratic Formula unless you don't see it the better way! You should recognize a Perfect Square Trinomial. Just grab these, write down the answer, and move on. \(x^{2} - 16x + 64 = (x-8)^{2}\) Rather suggests to me that we are not quite looking at the right problem statement.

OpenStudy (anonymous):

or apply completing the square....

OpenStudy (anonymous):

... got your point @tkhunny ...

OpenStudy (anonymous):

what about Given the points: A(0,0), B (6,3), C(5,5), and D(-1,2). What two segments must be parallel? Choose all possible answers. AB || DC AB || BC AB || AD DC || AD BC || AD BC || CD

OpenStudy (anonymous):

does anyone know the answer for that

OpenStudy (tkhunny):

@ShreeShree1501 Please post your own thread.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

to make it more easy use that method if just x^2 (a = 1) make ( x+)(x+) the last value it the product of the last two 64 = 8 *8 and the last sign tells if they have the same sign or not and the middle sign is for the largest value.

OpenStudy (anonymous):

I think I posted the wrong answers up there... These were the answers I meant to put: (x + 4)(x + 16) (x - 4)(x - 16) (x + 8)(x + 8) (x - 8)(x - 8)

OpenStudy (anonymous):

What would the answer be?

OpenStudy (tkhunny):

I sincerely hope no one will answer this last query. 1) You have shown no work of your own. None whatsoever. 2) You already have been given the answer. You should be here to learn, not to get someone else to do your homework for you. You tell me, with confidence, what the answer is.